A bathtub is being filled with water. After 3 minutes, 4/5 of the tub is full. Assuming the rate is constant, how much longer will it take to fill the bathtub?
My choices are a. 36 seconds b. 45 seconds c. 186 seconds d. 216 seconds

Answer by ptaylor(2198)

Let t=time required to finish filling the tub
The tub is being filled at the rate of (4/5)(1/3)=4/15 of the tub per min
After three minutes, the tub has filled (4/15)*3=4/5 of the tub leaving 1/5 of the tub yet to be filled
Soooo, our equation to solve is:
(4/15)*t=1/5
t=(1/5)/(4/15) multiply numerator and denominator by (15/4) to get rid of the complex fraction and we get
t=(1/5)(15/4)/1=15/20=(3/4) min=45 sec
Hope this helps---ptaylor